# Eccentricity signatures in LIGO-Virgo-KAGRA’s BNS and NSBH binaries

### v1

@@ IV.1 Parameter Estimation Challenges
We encountered several challenges when using the TEOBResumS-Dali and SEOBNRv5EHM waveform models, most notably the presence of complex multi-modal structure in the likelihood. These arise, for example, when varying the relativistic anomaly in SEOBNRv5EHM or the mean anomaly in TEOBResumS-Dali, and are also evident along mass parameters. Crucially, these features are not artifacts of under-sampling, but originate from the waveform models themselves. This can be demonstrated by computing the match between a fixed reference waveform and a set of waveforms with identical parameters except for eccentricity. The resulting match curve shows clear, repeated peaks corresponding to the distinct modes observed in Fig. 4. This strongly suggests that the features are imprinted directly in the waveform as a function of eccentricity or anomalies. The physical origin of these structures remains unclear and presents an interesting direction for future study. These complexities also complicate the inference process, as illustrated in Fig. 4, where even with most parameters fixed, the posterior remains multi-modal and will be further complicated when the full parameter space is explored.
@@ V Conclusion and Discussion
In this work, we conducted the first comprehensive analysis of eccentricity in six confidently detected low-mass gravitational-wave events reported by the LVK collaboration to date using a newly developed waveform model SEOBNRv5EHM. To assess the impact of waveform systematics, we cross-validate our results using the TEOBResumS-Dali model.
@@
Our findings reinforce the expectation that most low-mass binaries observed to date are consistent with circular inspiral (see Table 1). We also placed the first upper bound on the orbital eccentricity to be ranging from $e<0.063$ to $e<0.075$ for GW190814 and found no support for eccentricity in this event. Although, the inclusion of higher-order modes slightly reduces the eccentricity estimates in GW190814, the results remain broadly consistent across waveform models. In contrast, GW200105, show moderate evidence for non-zero eccentricity, with constraints of $e<0.14$ with 90% confident interval when eccentricity is sampled uniformly between 0 and 0.2. This result persists across both waveform models and inference settings, and remains consistent under a range of validation tests, including variations in the frequency cutoff, sampling diagnostics, and waveform tapering procedures. However, we find that the eccentricity constraint is sensitive to the choice of prior, as also noted in Refs. [15, 16]. When repeating the analysis using a prior uniform in $\log_{10}(e)$ with a lower bound of $10^{-4}$, we found that the support for moderate eccentricity is significantly reduced to twice in favor of eccentricity. This demonstrates that prior assumptions can significantly influence the inferred eccentricity distribution, particularly when the likelihood surface is broad or multi-modal.
@@
We ensure that the observed eccentricity in GW200105 is not an artifact of detector noise, waveform systematics, or prior assumptions by performing several verification tests. First, we visually inspected the data for glitches or signs of non-stationary Gaussian noise and found no anomalies that could mimic eccentricity. We then verified that our eccentric waveform models produced stable and well-behaved signals at higher eccentricities, with no signs of tapering issues or boundary artifacts. To test the robustness of the inference, we varied the low-frequency cutoff from 18 to 21 Hz while keeping the reference frequency fixed. The recovered eccentricity remained non-zero across these configurations and increased slightly at lower cutoffs, consistent with improved sensitivity to early inspiral features and no lower noise effects were present. Finally, we tested the loss in eccentricity information by inferring the eccentricity using 30 Hz as the starting frequency of our likelihood calculation and a reference and lower frequency of 20 Hz for the waveform generation. The eccentricity remained relatively high and we verified this result by generating an eccentric injection at the maximum-likelihood parameters from the original 30 Hz test. Our injection recovery was consistent with the eccentricity measured in the test. These cross-checks affirm that the measured eccentricity in GW200105 is not a numerical or modeling artifact, but a robust feature supported by multiple lines of evidence.
@@
Despite these findings, several factors highlight the need for continued investigation into the nature of eccentricity in compact binary mergers. One concern is the potential degeneracy between eccentricity and spin-induced precession. Although, this effect is less significant for low-mass systems like GW200105 [10, 15], it demands further study with improved waveform models and high signal-to-noise ratio events. Additionally, GW200105 was observed by only one detector, which limits our ability to fully rule out noise-induced features that that may mimic eccentricity. Nevertheless, our injection studies performed with a starting frequency of 30 Hz, recover eccentricities consistent with the injected values, making it unlikely that noise alone accounts for the observed signal. Finally, we also identified a secondary mode or island in the posterior distribution which involves a trade-off between the chirp mass and eccentricity, consistent with the findings in Morras et al. [15]. A bimodality near $e\sim 0.12$ in the main peak, previously reported by de Lluc Planas et al. [16], is also recovered in our analysis. Interestingly, while Morras et al. [15] and de Lluc Planas et al. [16] report slight disagreements between their two runs regarding the number and locations of posterior peaks, our analysis recovers all of these features with both waveform models. This suggests that, when considered together, the two prior studies are mutually consistent and that our results are compatible with both. The presence of these degeneracies underscores the complexity of parameter estimation in such systems, particularly when multiple physical effects persist.
@@
As the number of NSBH detections increases, so does the potential to verify formation channels that produce eccentric NSBH populations. The fourth observing run of the LIGO–Virgo–KAGRA collaboration (O4) has so far yielded approximately 300 low-latency alerts, including several NSBH candidates. With continued observations, these formation channels may soon be confirmed [87, 88]. As the current detector network grows in sensitivity and detection rate, next-generation observatories such as Cosmic Explorer [89, 90] and the Einstein Telescope [91, 92] are expected to observe thousands of mergers annually. This rapid growth will place increased demands on parameter estimation pipelines, particularly when using waveform models that include features like eccentricity, spin precession, and higher-order modes. These models are computationally expensive and require dense sampling of multi-modal parameter spaces in parameters such as mass, eccentricity, and anomaly.
@@
To make full use of the data provided by future detectors, more efficient inference techniques and waveform models are essential. In our analyses, the dominant computational cost arose from waveform generation. Reducing the waveform generation time is therefore a key priority. ROMs offer a promising route to accelerate models such as TEOBResumS-Dali and SEOBNRv5EHM, making them a critical focus for future development.
@@
Another key strategy for computational efficiency involves marginalization schemes that reduce sampling complexity by integrating over nuisance parameters. For instance, marginalizing over phase and anomaly could help mitigate multi-modality in the likelihood. Inference runs that included higher-order modes were approximately three times slower than those restricted to the dominant mode, primarily due to the inability to marginalize over the coalescence phase in comparison to our (2,2) mode runs. Future developments in marginalization methods could improve performance, especially for low-mass systems, where these modes are more influential.
@@
In this work, we have already experimented with marginalization over total mass to avoid sampling over the modes, and help reduce the number of waveform generations. The model we implemented takes an initial waveform and stretches or squeezes the waveform to the desired final total mass. This can help avoid having to generate multiple waveforms for a set of given parameters. We have not tested the model for the longest duration signals, but it may be beneficial towards BNS systems where waveform generation takes 10$\geq$ 10≥ 10 seconds to generate. In practice, we found the method to be inefficient for GW200105 and chose not to apply it in our final analyses.
@@
Understanding the degeneracy structure between anomaly, eccentricity, and mass could also lead to reparameterization using a pseudo-anomaly or effective eccentricity. Such transformatoin may help smooth over the discrete modes in the likelihood, effectively turning the distribution Gaussian and improving sampling efficiency. If the structured bands observed in waveform matches reflect unresolved modes, then designing samplers that can resolve these narrow likelihood features efficiently will be crucial. These advances will be necessary to enable fast and robust eccentricity measurements in the high-throughput era of the Cosmic Explorer and the Einstein Telescope and the anticipation of the O4 data release.
@@
The necessary information to reproduce these results and the posterior files can be found https://github.com/kkacanja/ecc_\__pe.git.
@@ Acknowledgements.
KK, KS, and AHN acknowledge support from the NSF grant PHY-2309240. The authors acknowledge the support from Syracuse University for providing computational resources through the OrangeGrid High Throughput Computing (HTC) cluster supported by the NSF award ACI-1341006. The authors express their sincere gratitude to Rossella Gamba for her invaluable assistance and for providing access to TEOBresumS, which greatly facilitated this work. KK also extends thanks to Yifan Wang for developing the plugin used to generate the gravitational waveforms in PyCBC and for the many helpful discussions, comments, and suggestions that contributed to this project.. Finally, we are deeply grateful to Aldo Gamboa and Alessandra Buonanno for permitting the use of the SEOBNRv5EHM eccentric waveform model, which was integral to our analysis, and for the feedback received on the paper. Additionally, KK would like to thank Aldo Gamboa for all the discussions related to the modes discovered in the eccentric waveforms. KK would like to thank Gonzalo Morras for allowing us to use their posteriors in our plots, and Maria de Lluc Planas et al. for making their posteriors publicly available. KK, KS, and AHN thank Jenny Vo and Paw Eh Blut Say, high school students who initially performed a preliminarily analysis to investigate the eccentricity in GW190814 through the Syracuse University Research in Physics (SURPh) program. Their early contributions helped lay the groundwork for what evolved into a broader investigation. This research has used data or software obtained from the Gravitational Wave Open Science Center (gwosc.org), a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.

### v2

@@ IV.1 Parameter Estimation Challenges
We encountered several challenges when using the SEOBNRv5EHM and TEOBResumS-Dali waveform models, most notably the presence of a complex multi-modal structure in the likelihood. These arise, for example, when varying the relativistic anomaly in SEOBNRv5EHM or the true anomaly in TEOBResumS-Dali, and are also evident along mass parameters. Crucially, these features are not artifacts of undersampling but rather originate from the waveform models themselves. This can be demonstrated by computing the match between a fixed reference waveform and a set of waveforms with identical parameters except for eccentricity. The resulting match curve shows clear repeated peaks corresponding to the distinct modes observed in Fig. 4. This strongly suggests that the features are imprinted directly in the waveform as a function of eccentricity or anomalies. The physical origin of these structures remains unclear and presents an interesting direction for future study. These complexities also complicate the inference process, as illustrated in Fig. 4, where even with most parameters fixed, the posterior remains multi-modal and will be further complicated when the full parameter space is explored.
@@ V Conclusion and Discussion
In this work, we conducted the first comprehensive analysis of eccentricity in seven confidently detected low-mass gravitational-wave events reported by the LIGO–Virgo–KAGRA (LVK) collaboration to date using a newly developed waveform model SEOBNRv5EHM. To assess the impact of waveform systematics, we cross-validate our results using the TEOBResumS-Dali model.
@@
Our findings reinforce the expectation that most low-mass binaries observed to date are consistent with circular inspirals (see Table 1). We also placed the first upper bound on the orbital eccentricity, ranging from $e<0.063$ to $e<0.075$ for GW190814, and found no support for eccentricity in this event. Although the inclusion of HOMs slightly reduces the eccentricity estimates in GW190814, the results remain broadly consistent across waveform models. In contrast, GW200105 shows moderate evidence for non-zero eccentricity, with constraints of $e<0.14$ at a 90% confidence interval when eccentricity is sampled uniformly between 0 and 0.2. This result persists across both waveform models and inference settings and remains consistent under a range of validation tests, including variations in the frequency cutoff, sampling diagnostics, and waveform tapering procedures. However, we find that the eccentricity constraint is sensitive to the choice of prior, as also noted in Refs. [19, 20]. When repeating the analysis using a prior uniform in $\log_{10}(e)$ with a lower bound of $10^{-4}$, we found that the support for moderate eccentricity is significantly reduced to twice in favor of eccentricity. This demonstrates that prior assumptions can significantly influence the inferred eccentricity distribution, particularly when the likelihood surface is broad or multi-modal.
@@
We ensure that the observed eccentricity in GW200105 is not an artifact of detector noise, waveform systematics, or prior assumptions by performing several verification tests. First, we visually inspected the data for glitches or signs of non-stationary Gaussian noise and found no anomalies that could mimic eccentricity. We then verified that our eccentric waveform models produced stable and well-behaved signals at higher eccentricities, with no signs of tapering issues or boundary artifacts. To test the robustness of the inference, we varied the low-frequency cutoff from 18 to 21 Hz while keeping the reference frequency fixed. The recovered eccentricity remained non-zero across these configurations and increased slightly at lower cutoffs, consistent with improved sensitivity to early inspiral features, and no lower noise effects were present. Finally, we tested the loss in eccentricity information by inferring the eccentricity using 30 Hz as the starting frequency of our likelihood calculation and a reference lower frequency of 20 Hz for waveform generation. The eccentricity remained relatively high, and we verified this result by generating an eccentric injection at the maximum-likelihood parameters from the original 30 Hz test. Our injection recovery was consistent with the eccentricity measured in the test. These cross-checks affirm that the measured eccentricity in GW200105 is not a numerical or modeling artifact, but a robust feature supported by multiple lines of evidence.
@@
Despite these findings, several factors highlight the need for continued investigation into the nature of eccentricity in compact binary mergers. One concern is the potential degeneracy between eccentricity and spin-induced precession. Although this effect is less significant for low-mass systems like GW200105 [10, 19], it demands further study with improved waveform models and high signal-to-noise ratio events. Additionally, GW200105 was observed by only one detector, which limits our ability to fully rule out noise-induced features that may mimic eccentricity. Nevertheless, our injection studies performed with a starting frequency of 30 Hz recover eccentricities consistent with the injected values, making it unlikely that noise alone accounts for the observed signal. Finally, we also identified a secondary mode or island in the posterior distribution which involves a trade-off between the chirp mass and eccentricity, consistent with the findings in Morras et al. [19]. A bimodality near $e\sim 0.12$ in the main peak, previously reported by de Lluc Planas et al. [20], is also recovered in our analysis. Interestingly, while Morras et al. [19] and de Lluc Planas et al. [20] report slight disagreements between their two runs regarding the number and locations of posterior peaks, our analysis recovers all of these features with both waveform models. This suggests that, when considered together, the two prior studies are mutually consistent and that our results are compatible with both. The presence of these degeneracies underscores the complexity of parameter estimation in such systems, particularly when multiple physical effects persist.
@@
As the number of NSBH detections increases, so does the potential to verify formation channels that produce eccentric NSBH populations. The fourth observing run of the LVK collaboration has so far yielded approximately 300 low-latency alerts, including several NSBH candidates. With continued observations, these formation channels may soon be confirmed [97, 98]. As the current detector network grows in sensitivity and detection rate, next-generation observatories such as Cosmic Explorer [99, 100] and the Einstein Telescope [101, 102] are expected to observe thousands of mergers annually. This rapid growth will place increased demands on parameter estimation pipelines, particularly when using waveform models that include features like eccentricity, spin precession, and higher-order modes. These models are computationally expensive and require dense sampling of multi-modal parameter spaces in parameters such as mass, eccentricity, and anomaly.
@@
To make full use of the data provided by future detectors, more efficient inference techniques and waveform models are essential. In our analyses, the dominant computational cost arose from waveform generation. Reducing the waveform generation time is therefore a key priority. ROMs offer a promising route to accelerate models such as SEOBNRv5EHM and TEOBResumS-Dali, making them a critical focus for future development.
@@
Another key strategy for computational efficiency involves marginalization schemes that reduce sampling complexity by integrating over nuisance parameters. For instance, marginalizing over phase and anomaly could help mitigate multi-modality in the likelihood. Inference runs that included higher-order modes were approximately three times slower than those restricted to the dominant mode, primarily due to the inability to marginalize over the coalescence phase compared to our (2,2) mode runs. Future developments in marginalization methods, such as the phase marginalization for higher order harmonics demonstrated in Roulet et al. [76], could improve performance, especially for low-mass systems, where these modes are more influential.
@@
In this work, we have already experimented with marginalization over the total mass to avoid sampling over the modes and help reduce the number of waveform generations. The model we implemented takes an initial waveform and stretches or squeezes it to achieve the desired final total mass. This can help avoid having to generate multiple waveforms for a set of given parameters. We have not tested the model for the longest duration signals, but it may be beneficial towards BNS systems where waveform generation takes $\geq 10$ seconds. In practice, we found the method to be inefficient for GW200105 and chose not to apply it in our final analyses.
@@
Understanding the degeneracy structure between anomaly, eccentricity, and mass could also lead to reparameterization using a pseudo-anomaly or effective eccentricity. Such a transformation may help smooth over the discrete modes in the likelihood, effectively turning the distribution Gaussian and improving sampling efficiency. If the structured bands observed in waveform matches reflect unresolved modes, then designing samplers that can efficiently resolve these narrow likelihood features will be crucial. These advances will be necessary to enable fast and robust eccentricity measurements in the high-throughput era of the Cosmic Explorer and the Einstein Telescope, and the anticipation of the O4 data release.
@@
The necessary information to reproduce these results and the posterior files can be found on GitHub [103]
@@ Acknowledgements.
KK, KS, and AHN acknowledge support from the NSF grant PHY-2309240. The authors acknowledge the support from Syracuse University for providing computational resources through the OrangeGrid High Throughput Computing (HTC) cluster supported by the NSF award ACI-1341006. The authors express their sincere gratitude to Rossella Gamba for her invaluable assistance and for providing access to TEOBresumS, which greatly facilitated this work. KK also extends thanks to Yifan Wang for developing the plugin used to generate the gravitational waveforms in PyCBC and for the many helpful discussions, comments, and suggestions that contributed to this project. Finally, we are deeply grateful to Aldo Gamboa and Alessandra Buonanno for permitting the use of the SEOBNRv5EHM eccentric waveform model, which was integral to our analysis, and for the feedback received on the paper. Additionally, KK would like to thank Aldo Gamboa for all the discussions related to the modes discovered in the eccentric waveforms. KK would like to thank Gonzalo Morras for allowing us to use their posteriors in our plots, and Maria de Lluc Planas et al. for making their posteriors publicly available. The authors thank Jenny Vo and Paw Eh Blut Say, high school students who initially performed a preliminary analysis to investigate the eccentricity in GW190814 through the Syracuse University Research in Physics (SURPh) program. Their early contributions helped lay the groundwork for what evolved into a broader investigation. This research has used data or software obtained from the Gravitational Wave Open Science Center [68, 69, 70], a service of the LIGO Scientific Collaboration, the Virgo Collaboration, and KAGRA. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation, as well as the Science and Technology Facilities Council (STFC) of the United Kingdom, the Max-Planck-Society (MPS), and the State of Niedersachsen/Germany for support of the construction of Advanced LIGO and construction and operation of the GEO600 detector. Additional support for Advanced LIGO was provided by the Australian Research Council. Virgo is funded, through the European Gravitational Observatory (EGO), by the French Centre National de Recherche Scientifique (CNRS), the Italian Istituto Nazionale di Fisica Nucleare (INFN) and the Dutch Nikhef, with contributions by institutions from Belgium, Germany, Greece, Hungary, Ireland, Japan, Monaco, Poland, Portugal, Spain. KAGRA is supported by Ministry of Education, Culture, Sports, Science and Technology (MEXT), Japan Society for the Promotion of Science (JSPS) in Japan; National Research Foundation (NRF) and Ministry of Science and ICT (MSIT) in Korea; Academia Sinica (AS) and National Science and Technology Council (NSTC) in Taiwan.